Presentation is loading. Please wait.

Presentation is loading. Please wait.

DØ FPD Overview Andrew Brandt UTA Q4 D S Q3S A1A2 P 1 UP p p Z(m) D1 Detector Bellows Roman Pot 233359 33230 57 P 2 OUT Q2 P 1 DN P 2 IN D2 Q4Q3Q2 FPD.

Similar presentations

Presentation on theme: "DØ FPD Overview Andrew Brandt UTA Q4 D S Q3S A1A2 P 1 UP p p Z(m) D1 Detector Bellows Roman Pot 233359 33230 57 P 2 OUT Q2 P 1 DN P 2 IN D2 Q4Q3Q2 FPD."— Presentation transcript:

1 DØ FPD Overview Andrew Brandt UTA Q4 D S Q3S A1A2 P 1 UP p p Z(m) D1 Detector Bellows Roman Pot 233359 33230 57 P 2 OUT Q2 P 1 DN P 2 IN D2 Q4Q3Q2 FPD Workshop February 20, 2003 Fermilaab

2 Workshop Goals Bring FPD detector group, proton-ID, and physics group together (if video allows), generate discussion Evaluate current status, problem areas, and issues Identify manpower shortages and attempt to address them Identify obstacles to producing physics results Develop plan for next six months

3 Diffraction Thesis Topics Soft Diffraction and Elastic Scattering: Inclusive Single Diffraction Elastic scattering (t dependence) Jorge Molina Total Cross Section Centauro Search Inclusive double pomeron Search for glueballs/exotics Hard Diffraction: Diffractive jet Michael Strang Diffractive b,c Diffractive W/Z Diffractive photon Diffractive top Diffractive Higgs Other hard diffractive topics Double Pomeron + jets Tamsin Edwards Other Hard Double Pomeron topics Rapidity Gaps: Central gaps+jets Topics in RED were studied Gap tags vs. proton tags with gaps only in Run I Double pomeron with gaps   E 1000 tagged events in Run II

4 All 6 castles with 18 Roman pots comprising the FPD were constructed in Brazil, installed in the Tevatron in fall of 2000, and have been functioning as designed. A2U pot had vacuum leak and was disabled for 6 months and fixed during Jan. 2003 shutdown. A2 Quadrupole castle installed in the beam line. Castle Status

5 Detector Setup 0.8 mm 3.2 mm 1 mm 17.39 mm U U’ X X’ V V’ Trigger 6 planes per detector in 3 frames plus a trigger scintillator U and V at 45 degrees to X, 90 degrees to each other Planes in a frame offset by ~2/3 fiber (4 fibers/bin) 2 detectors per spectrometer, 7 MAPMT per detector

6 At the University of Texas, Arlington, scintillating and optical fibers were spliced and inserted into the detector frames. Detector Assembly 20 detectors (includes 2 spares) built at UTA, sent to Fermilab, 10 installed, others in preparation at Fermi (see Mike S. talk).

7 All 18 cartridges have been assembled, 16 are installed in tunnel (10 with full detectors 2 with trigger scint). The 10 instrumented pots (Phase I) are ups, downs, and dipoles. Cables and tunnel electronics (low voltage, amp/shapers, etc.) installed and completely operational for Phase I, mostly operational for Phase II 18 pots setup (more later). 10 more detectors (includes 2 spares) are complete except for final polishing, which is in progress at Fermilab. Tunnel and Detector Status P2 Quadrupole castle with up and down detectors installed.

8 In the October 2001 shutdown four veto counters (designed at UTA, built at Fermilab) each of which cover 5.2 < |  | < 5.9 were installed between DØ and the first low beta quadrupole (Q4), about 6 m from the interaction point. The counters, two each on the outgoing proton and anti-proton arms, can be used in Diffractive triggering (veto proton remnant). Veto Counters

9 Pot Motion Pot motion is controlled by an FPD shifter in the DØ Control Room via a Python program that uses the DØ online system to send commands to the step motors in the tunnel.

10 Pot Motion Safeguards The software is reliable and has been tested extensively. It has many safeguards to protect against accidental insertion of the pots into the beam. The drivers are disabled with a switch in the Control Room when the pots are not being moved. The pots are hooked to an emergency line which bypasses the software to send the pots back to the home position in case of emergency.

11 Pot Insertion Monitor Effect of the pot motion on the proton and antiproton losses at DØ and CDF is monitored using ACNET. Current agreement with Beams Division and CDF requires that the effect on halo rates is less than 10%.

12 Recent FPD Milestones

13 FPD Contributors Brazil (LAFEX, UERJ, UFRJ, UNESP) UTA Los Andes (people) Fermilab($) Manchester(people) MSU ($) Nebraska ($) Thanks Greg! NIKHEF ($) NIU (Machine shop) Northwestern($) Notre Dame ($) Prague (people) Saclay (people) Manchester(people) Fermilab($) Can we generate more interest/support?

14 People+Jobs at Fermilab Leaving: Luiz, Jorge B. THANKS Jorge Leaving FPD: Victor THANKS Victor New people: Tamsin Edwards (grad student) Mary Elena Pol (UERJ faculty 2 mos) Continuing Full-time people: Mike S., Wanda, Carley, Ricardo Continuing Part-time people: Mike M., Vladimir FPD Experts: Carley, Wanda, Ricardo, Mike, Victor Trainees: Tamsin +? Jobs 1) Trigger integration and commissioning (all) 2) Expert shifts 3) Access coordination (?+AB) +accesses (all+?) 4) Run plan+shift schedule (MS+AB) 5) Friday 2 pm meeting (MS) 6) Examine (MS) 7) Database + Pot motion (WP+VS) 8) SCR triggering (MS?) 9) Data analysis (MS+TE) 10) Halo studies (MS+AB+WP+CM) *NEED MORE PEOPLE!

15 FPD January Shutdown Access Coordinator: Victor Bodyagin TUNNEL: 1)A2 vacuum repairs, requires coordination with BD and survey  (completed without bakeout) 2)Cable repairs  (only a few dead channels remain in phase I: note should produce list) 3)Detector installation (12 detectors and 6 pseudodetectors) ~  (10+6) 4)LVDT debugging  (bad LVDT found) 5)Pot motion upgrades ~ some done (see W.P talk) 6)Replacement of extension cables+ORC (not done) 7)Add some sandbags? (soon) COLLISION HALL: 1)PW08 preparation: installation of 3 TPP’s and AFE’s ~ 3 TPP’s installed, but only 2 AFE’s due to contact or noise problems (see V.B talk)

16 Access List TUNNEL: 1)Detector installation A1I+P1O to SCR (needs rearrangement of A/S to only use 11 boards) 2)Add sandbags at P2 and in front of electronics 3)Open air flow on P side 4)A2I LVDT replacement 5)A1O pot motion debugging 6)LMB pin/diode adapters 7)Phase II work (July shutdown) COLLISION HALL: 1)Debug AFE/TPP noise

17 FPD Operations FPD Operations Continue with expert shifters to insert pots and calorimeter shifters to monitor system and remove pots Re-establish operating positions, since may have changed due to access and A2 vacuum repairs Begin commisioning integrated FPD -Take calibration runs, understand noise -Finalize AFE timing -Take data recording FPD for all events Add new AND/OR terms and FPD triggers Work towards automated pot insertion Combine shifts with CFT (more in Mike S. talk)

18 AFE+DFE Integration Most of these items will be discussed in other talks— would like to end up with detailed work list + schedule Key:DONE ISSUES Know how to do I) AFE 1) Adding AFE, Sequencer, and VRB to database (VB+WP) 2) Pile-up tests for TPP (CA+JA) 3) Update FPD_AFE unpacking code (GA) 4) FPD_AFE examine (CA,VB,MS with Abid) 5) AFE+Seq Installation (GG, AFE’s temporarily) 5a) FIX SEQUENCER!! (RR with JA+FB 6) Method for installing 2 nd +3rd TPP's +installation (CA+RR+VB+Russ) 7) Update of AFE firmware for 3rd time (RR with JA,approval) 8) Start commissioning AFE (VB,MS,WP,CM II) DFE 1) Purchase of xilinx chips for first two boards (AB) 2) Construction of 2 DFE board (Jamieson) 3) Testing of DFE boards (RR+MV) 4) DFE firmware1 st draft (RR,MV,WC) 5) fixing DFE speed problem (RR+MV) 6) Testing of DFE boards+firmware in CTS (RR+MV) 7) Installation of 1 st DFE+LVDS cables (RR+MV) 8) DFE Examine/DFEware (RR) 9) Signals through to TM

19 Trigger Integration III) Luminosity Monitor/Trigger scintillator 1) LM TDC boards 2) LM Virtex boards 3) FPD_LM TDC boards (CM) 4) FPD_LM Virtex boards (CM) 5) LM Firmware (RR+MV?) 6) LM rates in EPICS (VB+V.Sirotenko) 7) FPD trigger from SCR until LM ready (VB) IV) Trigger Managaer 1) Installation of TM components Rob Mccrosky+CM 2) Cables connecting TM+3rd floor VB+Tom Diehl 3) First FPD multiplicity triggers 4) TM Firmware V) Phase II 1) 3 more TPP boards (4 th +5 th ) +spasre cut+stuffed (RR with Bob Jones) 2) Test new TPP boards (RR) 3) Install 4 th +5th TPP boards+AFE boards 4) Purchase 3 rd DFE board chips+install 5) Update TM firmware

20 Goals for 2003 Data taking with integrated Phase I Add FPD triggers to global list Complete funding and implementation of Phase II Preliminary results on several physics topics Would like to produce a detailed list in all areas: detector, software, PID, physics

21 Steps Needed for Physics Partial list of steps needed to get to physics quality data: Commission AFE’s Include trigger scintillator information/ basic and/or terms Alignment of FPD FPD database Activate FPD_Reco Halo understanding, reduction, and rejection FPD info in DST’s and Thumbnails Propose some reorganization to address some of these issues. 1)Alignment Task Force to address alignment, calibration, folding in accelerator information 2)Halo Task Force to propose studies for reducing halo and maximizing acceptance

22 How to apply accelerator information to improve beam position measurement A/S pedestal fluctuations, need to evaluate level of problem, develop hot and dead channel list Decoupling LMB from calorimeter …what is on your list? Random FPD Issues

23 Vacuum Leak Pilot Error Controlled Lines On Controlled Lines Bypassed Controlled Access Smoke Alarms Helium Limit Switches Sequencer TPP Discussion of Items from FPD-Probs Folder

24 Preparation forPhase II Requirements: About $90,000 primarily for MAPMT for 6 more detectors, test MAPMT, install new detectors Upgraded 12V power supplies to accommodate 14 A/S boards per quadrupole castle (A/S boards are in place) 2 more AFE/TPP boards (TPP boards and components exist but not yet stuffed), solve mechanical installation issues One more DFE board (need to purchase chips, stuff board) Some short cables from detector to A/S need to be repaired and tested Replace A2I LVDT Debug A1O IIB Purchase spares for critical systems Any cable issues?

25 MRI for Phase II ItemCost Graduate Student$49,000 Equipment$92,000 Indirect Cost$36,000 Miscellaneous$2,000 Total$179,000 NSF$125,000 Match$54,000 EquipmentCost MAPMT (56)$73,000 PMT (10)$6,400 LVDT (3)$2,200 Drivers/motors(3)$2,800 Xilinx chips$2,500 LV power$4,600 Total$92,000 Submitted with NIU, should hear about mid-summer (50%?)

26 Elastic Data t distribution Before alignment After alignment Gaussian fit

27 Timing is Everything We identify bunch structure of both proton and antiproton beam We can reject proton halo at dipoles using TDC timing CDF does not have this capability D1 TDC D2 TDC pbar p

28 Lots of progress Lots still to do Lets get to the physics Conclusions

Download ppt "DØ FPD Overview Andrew Brandt UTA Q4 D S Q3S A1A2 P 1 UP p p Z(m) D1 Detector Bellows Roman Pot 233359 33230 57 P 2 OUT Q2 P 1 DN P 2 IN D2 Q4Q3Q2 FPD."

Similar presentations

Ads by Google